OpenStudy (anonymous):

integrate (pi)dx / sqrt (pi)x^2

OpenStudy (anonymous):

\[\frac{ \Pi dx }{ \sqrt{\Pi x^{2}} }\]

OpenStudy (anonymous):

note that since \(\pi/\sqrt{\pi}=\sqrt\pi\) is a constant it passes right thru:$$\int\frac{\pi}{\sqrt{\pi} x^2}dx=\int\frac{\sqrt\pi}{x^2}dx=\sqrt{\pi}\int\frac1{x^2}dx$$

OpenStudy (anonymous):

now observe \(1/x^2=x^{-2}\) and use the power rule: $$\int x^ndx=\frac1{n+1}x^{n+1}+C$$

OpenStudy (anonymous):

sqrt (pi x^2)

OpenStudy (anonymous):

x^2 is inside the sqrt. :) so x^2 becomes x

OpenStudy (ranga):

sqrt(pi) * ln(x) + C

OpenStudy (anonymous):

okay, not a worry: $$\sqrt{\pi x^2}=\sqrt{\pi}\sqrt{x^2}=\sqrt{\pi}|x|$$so the constant is the same but now we integrate \(1/|x|\) rather than \(1/x^2\). this is a little trickier. if you're told that \(x\ge0\) then this is just \(1/x\) of course

OpenStudy (anonymous):

$$\sqrt\pi\int\frac1{|x|}dx=\sqrt\pi\int\frac1xdx=\sqrt{\pi}\log |x|+C$$

OpenStudy (anonymous):

for \(x\ge 0\)

OpenStudy (ranga):

logarithm does not exist for negative numbers. Also, the original function has an x in the denominator and therefore x = 0 is not in the domain of the function. So sqrt(pi) * ln(x) + C for x > 0.

OpenStudy (anonymous):

just one question

OpenStudy (anonymous):

why is it many uses log and ln interchangably

OpenStudy (anonymous):

? :)

OpenStudy (ranga):

In calculus natural logarithm is what is meant when we say logarithm unless the problem explicitly says what the base is.

OpenStudy (anonymous):

natural logarithm is well natural in the sense it is the logarithmic function that gives the area bounded by the positive branch of \(1/x\)

OpenStudy (anonymous):

and anyways @ranga is correct that this particular function 1/x is not integrable over any non-trivial interval containing 0 but that's not in general true for other functions. take for example:$$\int_{-1}^2\frac1{x^{1/2}}dx$$even though \(1/x^{1/2}\) is undefined at \(x=0\) the function is still integrable over that piece. although, yes, you are correct that \(\log x\) is undefined at \(x=0\)

OpenStudy (anonymous):

engineers use \(\ln\) because they are filthy

OpenStudy (anonymous):

ok. thanks guyz!! :D

OpenStudy (anonymous):

just for the note, :) i'm an engineer. ;)

OpenStudy (ranga):

\[\frac{ d }{ dx}\ln(x) = \frac{ 1 }{ x }\quad but \quad \frac{ d }{ dx}\log_{10} (x) = \frac{ 1 }{ \ln10 } \times \frac{ 1 }{ x }\]

OpenStudy (anonymous):

i mean.. not yet an engineer but becoming an engineer. :)